OpenStudy (anonymous):

which of the following lines does not intersect the line y=-2x +3 A)y=2x B)y=x-3 C)y=2x+3 D)y=-2x-3

OpenStudy (haseeb96):

D is the correct answer because when u will put y =-2x-3 at the placce of y in the given equation then the equation would not be existed y=-2x+3 y=-2x-3 now the left hand side of both equation same so compare both the right hand sides equations -2x-3=-2x+3 -2x+2x-3-3=0 -6=0 so this is not satisfy or not exist so D is the corrrect answer this line will not intersect

OpenStudy (haseeb96):

have u understood it if u have any difficulty then ask me ? i will try my best to help u ....

OpenStudy (anonymous):

I understand